# Relevance for Human Robot Collaboration

### v2

@@ Abstract
Effective human-robot collaboration (HRC) requires the robots to possess human-like intelligence. Inspired by the human’s cognitive ability to selectively process and filter elements in complex environments, this paper introduces a novel concept and scene-understanding approach termed ‘relevance.’ It identifies relevant components in a scene. To accurately and efficiently quantify relevance, we developed an event-based framework that selectively triggers relevance determination, along with a probabilistic methodology built on a structured scene representation. Simulation results demonstrate that the relevance framework and methodology accurately predict the relevance of a general HRC setup, achieving a precision of 0.99 and a recall of 0.94. Relevance can be broadly applied to several areas in HRC to improve task planning time by 79.56% compared with pure planning for a cereal task, reduce perception latency by up to 26.53% for an object detector, improve HRC safety by up to 13.50% and reduce the number of inquiries for HRC by 80.84%. A real-world demonstration showcases the relevance framework’s ability to intelligently assist humans in everyday tasks.
@@ I Introduction
Robots and automated systems are increasingly integrated into modern society, and the human world, which requires them to develop human-like intelligence and cognitive abilities [1, 2]. One example of such abilities is that human beings have the capability and tendency to focus on the relevant physical elements in an environment and relevant abstract concepts or information in the input based on the context, the objectives, the historical experiences, the reasoning, etc [3]. This ability of humans is demonstrated in various studies [4, 5].
@@
Empowering robots with this human-like cognitive capability offers two significant benefits. First, robots interacting closely with human beings will achieve a scene-understanding similar to that of humans. This results in more natural and effective behavior and interaction. Second, by identifying the most relevant elements in the scene, the robot can effectively focus its limited computation power on the relevant elements to reduce the computation requirements, increase the computation speed, and even increase the safety of the robots in HRC tasks [3].
Previous works have primarily focused on visual saliency and attention mechanisms to identify prominent features. However, they often overlook the broader context necessary for long-term task planning and proactive assistance. Additionally, attention mechanisms are typically constrained to short-term simple motion generation and lack interpretability, which makes them unsuitable for complex problems requiring comprehensive reasoning. These limitations highlight the need for a novel approach designed for context-aware, efficient, and accurate scene understanding, reasoning about the interconnection among humans, dynamic environments, context, etc.
In this paper, we define and introduce a novel concept and approach for scene understanding, termed ‘relevance’. In relevance, our approach to scene understanding starts by identifying all elements in the scene, which are then grouped into classes. Our approach identifies the classes of contextual elements pertinent to what is or may happen in the scene. Our approach helps identify all classes with supplementary elements. Such classes/elements are not core to the scene with a positive (enhancing), negative (distracting), or neutral (optional) effect. Finally, the approach identifies the core or essential elements, focusing on indispensable components like main characters, primary actions, and key elements for what is taking place. This approach ensures a structured, coherent, systematic, and efficient scene understanding that is updated in real time.
@@
To quantify relevance, we propose and develop a novel framework. The event-based framework contains modules of perception, triggers, relevance determination, and decision-making. The event-based mechanism initializes the relevance determination only under certain circumstances, resulting in a dramatic increase in computation speed. In the relevance determination module, a unified probabilistic methodology based on a novel scene representation is proposed and developed. This methodology accurately quantifies the relevance of the scene, which dramatically improves the performance of HRC.
@@
To summarize, the contributions of this paper are four-fold: 1. We proposed a new concept and scene understanding approach called relevance to identify the relevant elements in the scene for the human’s objective. 2. We developed an event-based framework for faster relevance determination. 3. We developed a systematic methodology for relevance quantification and determination. 4. We demonstrated the effectiveness of our framework and methodology using both simulation evaluation and real-world demonstration.
@@ II Related Works
Two areas of research, visual saliency and attention, share similar goals to our work but are fundamentally different. To the best of our knowledge, this work is the first in HRC to empower robots with cognitive ability by identifying and focusing on the relevant elements based on long-term objectives and context.
@@ II-A Visual Saliency
Visual saliency seeks to identify and localize the most distinctive regions or objects within an image based on visual conspicuity without considering the context. Traditional methods are developed based on the features at the pixel level, including colors, intensity, orientation, etc [6, 7, 8]. Deep learning-based methods can also be applied [9, 10, 11, 12]. However, visual saliency only considers visual conspicuity, neglecting important considerations, such as context, for proactive assistance in HRC.
@@ II-B Attention
Attention is another mechanism that selectively focuses on relevant parts of the input data but is dramatically different from relevance. In a nutshell, attention is structured with self-attention and/or cross-attention to formulate and optimize weights for the input features [13]. However, attention in robotics is only applied to end-to-end algorithms for better completion of the current short-term action, e.g., pick and place [14, 15, 16]. Moreover, attention has several challenging limitations, such as limited applicability only to end-to-end methods, limited interpretability, and myopia. Relevance is a fundamental function in robot intelligence for high-level human-like reasoning by considering human factors (long-term objectives, goals, preference, and motion), task models (task structure, task prediction, object applicability), environmental factors, etc.
From another perspective, applying attention to human-robot collaboration (HRC) is challenging due to three factors: the proactive nature of HRC, its inherent complexity, and the difficulty of dataset collection. First, attention mechanisms are effective in reactive tasks where the neural networks focus on specific parts of the immediate input to generate actions. However, for proactive assistance in general HRC, the model is supposed to consider long-term goals, which require projection into the future. Second, the complexity of HRC is very high, where the robots are supposed to reason the interconnection between the context, the tasks in a hierarchy, and the human (objective and preferences), which is challenging for attention mechanisms to capture. Third, end-to-end learning, where attention works, requires intensive data collection, which is challenging because of the dynamics and the higher dimension of the environment involving human actions.
@@ III-A Perception module
The perception module continuously processes the sensor input (camera, microphone, etc.) and obtains the features $\mathcal{F}$ of the scene. Unlike other pipeline frameworks for robots that process the perception and decision-making in a serial manner [17], our developed and proposed framework perceives the scene in an uninterrupted manner.
@@ III-B Triggers Check
The triggers $T$ are to determine if a subprocess should be initialized for relevance determination based on the features $\mathcal{F}$. Some potential criteria for trigger activation include element changes and/or updated objectives or new tasks for the human.
@@ III-C Relevance Determination
Our relevance determination methodology is a probability framework based on a new type of scene representation, i.e. classes and elements. The relevance determination methodology is a two-step methodology that determines the relevance classes and then the relevance elements. At the same time, the methodology determines if sufficient information is available. With sufficient information, the framework will go through the methodology and determine relevance of the scene. With information insufficient, the algorithms will request or wait for the perception module to give out more information and cues. More details of our methodology for relevance determination will be discussed in Section IV.
@@ IV-B Relevance Determination Mechanism
There are two types of outcomes from the mechanism $\mathcal{M}$. Suppose the information is sufficient for the relevance quantification and determination. In this case, the outcomes are the relevance scores for each class or element, and, at the same time, the set of relevant classes $\mathcal{C}_{r}$ or the set of relevant elements $\mathcal{E}_{r}$. If the information in $\mathcal{F}$ is not sufficient for relevance quantification and determination, then the algorithm will refer to the perception module to either actively request or obtain new cues $\mathcal{F}_{new}$ from the scene before continuing the processing. Thus, the mechanism $\mathcal{M}$ can be mathematically represented as:
$\mathcal{M}:\mathcal{F},\mathcal{X}\rightarrow\left\{\begin{array}[]{ll}\mathcal{R}(x)\text{ for }\forall x\in\mathcal{X}\text{, and }\mathcal{X}_{r}&\text{if }\text{sufficient }\mathcal{F}\\ \text{ask or wait for }\mathcal{F}_{new},&\text{else}\end{array}\right.$ (2)
@@ IV-C Class Relevance Modeling
In (7), the first term $P(\mathcal{C}_{i}\mid\mathcal{T},\mathcal{O})$ represents the probability that class $C_{i}$ is relevant given the task is $\mathcal{T}$, and the objective is $\mathcal{O}$. The second term $P(\mathcal{T}\mid\mathcal{O})$ represents the probability that the current task is $\mathcal{T}$ given the objective is $\mathcal{O}$. If the human’s historical preference data about the specific class $C_{i}$ for the task $\mathcal{T}$ and objective $\mathcal{O}$ is available, then the first term can be derived as the probability of $C_{i}$ is relevant based on the preference. If the historical preference data is not available for class $C_{i}$ for task $\mathcal{T}$, the probability $P(C_{i}\mid\mathcal{T},\mathcal{O})$ can be derived from predictions based on Large Language Models (LLM), other datasets that incorporate action sequences, other human cues, etc. The second term in (7) can be derived from prediction models for human tasks.
@@ IV-D Element Relevance Modeling
where $e_{j}\in C_{i}\in C_{r}$, and the symbol $\neg$ denotes logical negation. $P(e_{j}|\neg C_{i},\mathcal{O})$ represents the probability that $e_{j}$ is relevant given the class $e_{j}$ belongs to is not relevant and the objective is $\mathcal{O}$, which equals to 0. Thus, (9) becomes:
@@ V-B Relevance Quantification Evaluation
In this paper, we emphasize the simulation for the coffee domain for the brevity and clarity of our analysis. In the simulation, the objective is predefined as “get something to drink at the break of a conference”, and the task is defined as “drink cold brew coffee”. A large language model (LLM) is adopted to classify all the elements into the scene representation $\mathcal{S}$ and provide the probabilities necessary for the relevance quantification. We tested 25 combinations of the class thresholds $\tau_{c}$ and element thresholds $\tau_{e}$, with 30 simple cases randomly generated for each combination.
The simulation results are shown in Fig. 2, which agree well with common sense. As $\tau_{c}$ and $\tau_{e}$ increase, more classes and elements will be pruned. Thus, the recall decreases, the precision increases, and the relevant object ratio decreases as these two thresholds increase. the F1 scores demonstrate a trend that increases first and decreases with the threshold increase.
@@
We select the thresholds $\tau_{c}$ and $\tau_{e}$ to be 0.2. The number of cases in which each type of element is relevant is shown in Fig. 3 without considering the constraints. Of the 30 cases, a couple randomly add an ice bucket to the setup and predict that the ice bucket is relevant. All 30 cases predict that coffee and plastic cups are relevant. These results further demonstrate that relevance prediction is very precise.
@@ V-C Relevance in HRC
One benefit of relevance in HRC is that relevance can help the robot better reason about the scene and the human’s requirements in a well-structured manner, which resembles how humans interact with other humans. This benefit can be demonstrated with a simple robot inquiry example in a coffee serving setup. The human-like robot will first detect the relevant classes and types of elements. If the relevance of any classes or types of elements is high, those classes or types are deemed necessary (coffee in this domain). Other relevant classes or types of elements are optional, and the robot asks the human if they need those relevant elements. The number of inquiries required using different methods is shown in Table I. It is shown that with precise and proper reasoning about the relevance of an element to a task or objective, the number of inquiries required for appropriate assistance is reduced by 80.84% compared with the number without relevance. Thus, human-robot interactions can be much more natural and fluent.
@@ V-D Relevance in Task Planning
This section demonstrates the effectiveness of relevance in task planning. A Fast Downward planner is employed. We compare three methods: our relevance with 0.2 as thresholds, random relevance with 0.2 as thresholds, and pure planning without relevance. For each method, we evaluate the processing time (relevance computation + planning), the timeout rate (120s), and the failure rate. Results are shown in Table II.
Relevance dramatically improves task planning performance. Compared with Pure Planning, relevance significantly reduces the time cost by up to 79.56% and the timeout rate across all problem formulations. As the problem formulation becomes more complex, we observe a notable increase in the timeout rate and planning time for Pure Plan, highlighting its inefficiency in solving complex task-planning problems. When compared with random relevance, we observe a substantial reduction in failure rates, demonstrating that our relevance successfully and accurately determines the relevant elements of the scene.

### v4

@@ Abstract
Inspired by the human ability to selectively focus on relevant information, this paper introduces relevance, a novel dimensionality reduction process for human-robot collaboration (HRC). Our approach incorporates a continuously operating perception module, evaluates cue sufficiency within the scene, and applies a flexible formulation and computation framework. To accurately and efficiently quantify relevance, we developed an event-based framework that maintains a continuous perception of the scene and selectively triggers relevance determination. Within this framework, we developed a probabilistic methodology, which considers various factors and is built on a novel structured scene representation. Simulation results demonstrate that the relevance framework and methodology accurately predict the relevance of a general HRC setup, achieving a precision of 0.99, a recall of 0.94, an F1 score of 0.96, and an object ratio of 0.94. Relevance can be broadly applied to several areas in HRC to accurately improve task planning time by 79.56% compared with pure planning for a cereal task, reduce perception latency by up to 26.53% for an object detector, improve HRC safety by up to 13.50% and reduce the number of inquiries for HRC by 80.84%. A real-world demonstration showcases the relevance framework’s ability to intelligently and seamlessly assist humans in everyday tasks.
@@ I Introduction
Robots and automated systems are increasingly integrated into modern society, and the human world, which requires them to develop human-like intelligence and cognitive abilities [1, 2]. One example of such abilities is the human capability to selectively focus on relevant physical elements in the environment and relevant abstract concepts or information in the input, guided by context, objectives, historical experiences, and reasoning. This cognitive function is closely associated with the reticular activating system (RAS), a network of neurons in the brain responsible for filtering sensory information, capturing relevant details, and suppressing irrelevant stimuli to optimize cognitive processing and decision-making [3]. Various studies have demonstrated the aforementioned human ability. [4, 5].
@@
Empowering robots with this human-like cognitive capability offers two significant benefits. First, robots interacting closely with human beings will achieve a scene-understanding similar to that of humans. This results in more natural and effective behavior and interaction [6]. Second, by identifying the most relevant elements in the scene, the robot can effectively focus its limited computation power on the relevant elements to reduce the computation requirements, increase the computation speed, and even increase the safety of the robots in HRC tasks [7].
In this paper, we define and introduce a novel concept and approach for dimensionality reduction, termed ‘relevance.’ Relevance is defined as a dimensionality reduction process that continuously perceives the scene, identifies and detects its elements, and organizes them into a hierarchical representation, such as attribute classes. This process utilizes contextual cues to selectively reduce input dimensionality. When cues are insufficient, relevance-driven processing iteratively gathers additional cues and reapplies dimensionality reduction until sufficient information is obtained.
Previous works in this area to identify prominent features have primarily focused on visual saliency and attention mechanisms. Visual saliency aims to identify the most distinctive regions within an image based on the image features, such as color, intensity, and orientation [8]. As a result, the salient regions detected are mostly visually conspicuous regions. However, visual saliency does not consider context information, which is critical in the application of HRC. Attention is a type of mechanism to build a mapping from the embedded input data to importance weights through a structured learnable function. The attention mechanism acts as an add-on module inside neural network models to consider the weights of each component to improve performance. Attention has been widely adopted and implemented in various applications, such as end-to-end frameworks to generate actions from sensory inputs [9], various sub-functions in the robotic pipelines [10, 11], etc.
Compared with saliency and attention, relevance is a comprehensive framework and concept that dramatically extends the scope and functionalities of these works. Our relevance, in its current form, already considers context information, such as human objectives, tasks, environment information, human preferences, etc. These factors are essential to enable proactive robotics assistance. Moreover, relevance is not only a module to generate the importance weight of each object in the input but also a novel framework that integrates a continuously running multi-modality perception module, a novel hierarchical scene representation, consideration of cue sufficiency in the input, etc. Those functions of relevance, which are not considered in saliency and attention, are essential for accurate, reliable, and seamless HRC. Last but not least, our relevance features a flexible formulation and computational principle not limited to the structure of neural networks. In this paper, as an illustration. we develop a probabilistic model that leverages an AI toolkit with a large language model (LLM), human preferences, low-level attributes, and constraints to generate the necessary information. This structure and formulation enable anticipatory capabilities, interactions among different information sources, a far more complex reasoning process, etc, which finally aims for artificial general intelligence (AGI). The demonstration in this paper shows, with relevance, that the robot can generate optimal reactions and seamless assistance to humans without transfer learning.
@@
To quantify relevance, we propose and develop a novel framework. The event-based framework contains modules of perception, triggers, relevance determination, and decision-making. The perception module is continuously running and perceives cues in the environment based on multi-modality sensors and processing. The event-based mechanism initializes the relevance determination only under certain circumstances, resulting in a dramatic increase in computation speed. In the relevance determination module, flexible formulations and computation principles can be applied. As an illustration, we developed a unified probabilistic methodology based on a novel scene representation. This methodology accurately quantifies the relevance of the scene, which dramatically improves the performance of HRC. Finally, a decision making module generates actions based on the results of relevance determination.
@@
To summarize, the contributions of this paper are four-fold: 1. We proposed a new concept, dimensionality reduction approach, and framework called relevance to enable fast, accurate, seamless, and proactive HRC. 2. We developed an event-based framework for faster relevance determination. 3. We developed a probabilistic methodology for relevance quantification and determination based on a novel hierarchical scene representation. 4. We demonstrated the effectiveness of our framework and methodology using both simulation evaluation and real-world demonstration.
@@ II Related Works
Two areas of research, i.e., visual saliency and attention, share similar goals to our work but with much narrower scope and functionalities. We also include a review of related HRC literature. To the best of our knowledge, this work is unique in HRC that it empowers the robots with the cognitive capability to focus on selective components in the input by developing a novel framework and methodology with unique advantages, which are essential for HRC.
@@ II-A Visual Saliency
Visual saliency seeks to identify and localize the most distinctive regions or objects within an image based on visual conspicuity without considering the context. Traditional methods are developed based on the features at the pixel level, including colors, intensity, orientation, etc [8, 12, 13]. Deep learning-based methods can also be developed and trained with the help of visual saliency datasets that are annotated either manually with a mouse click or automatically with eye gaze tracker [14, 15, 16, 17]. However, visual saliency mainly considers visual conspicuity, neglecting important considerations, such as context, for proactive assistance in HRC.
@@ II-B Attention
Attention is another mechanism that selectively focuses on relevant parts of the input data but with limited scope and functionalities compared to relevance. In a nutshell, attention is structured with self-attention, cross-attention, multi-layer perception (MLP), etc., to formulate and optimize weights for the input features [18]. However, attention in robotics is mainly applied to end-to-end algorithms for better completion of the short-term action, e.g., pick and place [9, 19, 20]. Moreover, attention has several challenging limitations, such as limited applicability to end-to-end methods only, limited interpretability, and myopia. Relevance not only resembles the cognitive capability to reduce the dimension of original input but also alleviates the limitations of attention by developing a novel event-based framework and distinctive methodologies, such as a flexible formulation, continuously running perception, hierarchical scene representation, multi-level relevance determination, etc. Thus, our relevance can integrate the information and cues from an AI toolkit with LLMs, human preferences, human objectives, environmental constraints, etc., to enable accurate, seamless, and proactive assistance in HRC.
II-C Human Robot Collaboration
There have been significant advancements in the fields related to HRC. We classify those works into two categories. The first category of works focuses on one specific important function for HRC, such as object detection [7], semantic segmentation [21], human intention prediction [22, 23, 24, 25], human motion prediction [26], human-aware task planning [27], etc. Those works, though important, have fundamentally different purposes from relevance. They are not for dimensionality reduction. The results of those modules can be applied for relevance quantification, and relevance can, vice versa, benefit these modules. The second category of work aims to develop a framework that directly generates actions from sensory inputs. Those works include learning from demonstration [28], reinforcement learning [29], etc. To incorporate the selective processing idea into these functions, some work leverage attention mechanisms if the model is neural network-based [30]. However, because of the narrower scope of attention, as described in Section II.B, our relevance possesses more functions and is more advanced than attention. As shown in Sections V and VI, with relevance, robots can generate proactive, seamless, accurate, and structured reactions in an autonomous manner.
@@ III-A Perception module
The perception module continuously processes the sensor inputs (camera, microphone, etc.) with an AI toolkit and obtains the features $\mathcal{F}$ of the scene. Unlike other pipeline frameworks for robots that process the perception and decision-making in a serial manner [26], our developed and proposed framework perceives the scene in an uninterrupted manner to capture the fast and dynamic changes in the scene. Potential algorithms in the AI toolkit include You Only Look Once (YOLO) [31], MMPose [32], CLIP [33], Visual Language Models (VLM) [34], etc.
@@ III-B Triggers Check
The triggers $T$ are to determine if a subprocess should be initialized for relevance determination based on the features $\mathcal{F}$ derived from the perception module. Some potential criteria for trigger activation include element changes and/or updated objectives or new tasks for the human.
@@ III-C Relevance Determination
One unique advantage of relevance is that it enables flexible formulations and computation principles. In this paper, as an illustration, our relevance determination methodology is a probability framework based on a new type of scene representation, i.e., classes and elements. The relevance determination methodology is a two-step methodology that determines the relevance classes and the relevance elements. At the same time, the methodology determines if sufficient information is available. With sufficient information, the framework will go through the methodology and determine relevance of the scene. With insufficient information, the algorithms will request or wait for the perception module to give out more information and cues. More details of our methodology for relevance determination will be discussed in Section IV.
@@ IV-B Relevance Determination Mechanism
There are two types of outcomes from the mechanism $\mathcal{M}$. Suppose the information is sufficient for the relevance quantification and determination. In this case, the outcomes are the relevance scores for each class or element, and, at the same time, the set of relevant classes $\mathcal{C}_{r}$ or the set of relevant elements $\mathcal{E}_{r}$. If the information in $\mathcal{F}$ is not sufficient for relevance quantification and determination, then the algorithm will refer to the perception module to request or obtain new cues $\mathcal{F}_{new}$ from the scene before continuing the processing. Thus, the mechanism $\mathcal{M}$ can be mathematically represented as:
$\mathcal{M}:\mathcal{F},\mathcal{X}\rightarrow\left\{\begin{array}[]{ll}\mathcal{R}(x)\text{ }\forall x\in\mathcal{X}\text{, and }\mathcal{X}_{r}&\text{if }\text{sufficient }\mathcal{F}\\ \text{ask or wait for }\mathcal{F}_{new},&\text{else}\end{array}\right.$ (2)
@@ IV-C Class Relevance Modeling
In (7), the first term $P(\mathcal{C}_{i}\mid\mathcal{T},\mathcal{O})$ represents the probability that class $C_{i}$ is relevant given the task is $\mathcal{T}$, and the objective is $\mathcal{O}$. The second term $P(\mathcal{T}\mid\mathcal{O})$ represents the probability that the current task is $\mathcal{T}$ given the objective is $\mathcal{O}$. If the human’s historical preference data about the specific class $C_{i}$ for the task $\mathcal{T}$ and objective $\mathcal{O}$ is available, then the first term can be derived as the probability of $C_{i}$ is relevant based on the preference. If the historical preference data is not available for class $C_{i}$ for task $\mathcal{T}$, the probability $P(C_{i}\mid\mathcal{T},\mathcal{O})$ can be derived from predictions based on the tools in the AI toolkit, such as Large Language Models (LLM), other datasets that incorporate action sequences, other human cues, etc. The second term in (7) can be derived from prediction models for human tasks.
@@ IV-D Element Relevance Modeling
where $e_{j}\in C_{i}\in C_{r}$, and the symbol $\neg$ denotes logical negation. $P(e_{j}|\neg C_{i},\mathcal{O})$ represents the probability that $e_{j}$ is relevant given the class that $e_{j}$ belongs to is not relevant and the objective is $\mathcal{O}$, which equals to 0. Thus, (9) becomes:
@@ V-B Relevance Quantification Evaluation
In this paper, we emphasize the simulation for the coffee domain for the brevity and clarity of our analysis. In the simulation, the objective is predefined as “get something to drink at the break of a conference”, and the task is defined as “drink cold brew coffee”. A large language model (LLM) GPT-4o is adopted to classify all the elements into the scene representation $\mathcal{S}$ and provide the probabilities necessary for the relevance quantification, as illustrated in Section IV. We tested 25 combinations of the class thresholds $\tau_{c}$ and element thresholds $\tau_{e}$, with 30 simple cases randomly generated for each combination.
The simulation results are shown in Fig. 2, which agree well with common sense. As $\tau_{c}$ and $\tau_{e}$ increase, more classes and elements will be pruned. Thus, the recall decreases, the precision increases, and the relevant object ratio decreases as these two thresholds increase. The F1 scores demonstrate a trend that increases first and then decreases with the threshold increase.
@@
We select the thresholds $\tau_{c}$ and $\tau_{e}$ to be 0.2. At this threshold combination, our methodology archives a recall $\mathfrak{R}$ of 0.94, a precision $\mathfrak{P}$ of 0.99, an F1 score $\mathfrak{F}$ of 0.96, and an object ratio $\mathfrak{N}$ of 0.94. These results demonstrate that our methodology is effective and accurate in dimension reduction. The number of cases in which each type of element is relevant is shown in Fig. 3 without considering the constraints. Of the 30 cases, a couple of them randomly add an ice bucket to the setup and predict that the ice bucket is relevant, which agrees with common sense in the context of drinking cold brew coffee. All 30 cases predict that coffee and plastic cups are relevant. These results further demonstrate that relevance prediction is very accurate.
@@ V-C Relevance in HRC
One benefit of relevance in HRC is that relevance can help the robot better reason about the scene and the human’s requirements in a well-structured manner, which resembles how humans interact with other humans. This benefit can be demonstrated with a simple robot inquiry example in a coffee serving setup. The human-like robot will first detect the relevant classes and types of elements. If the relevance of any classes or types of elements is high, those classes or types are deemed necessary (coffee in this domain). Other relevant classes or types of elements are optional, and the robot asks the human if they need those relevant elements. The number of inquiries required using different methods is shown in Table I. Note that the number of inquiries required without relevance equals the average number of objects in the testing setup. It is shown that with accurate and proper reasoning about the relevance of an element to a task or objective, the number of inquiries required for appropriate assistance is reduced by 80.84% compared with the number without relevance. Thus, human-robot interactions can be much more natural and fluent.
@@ V-D Relevance in Task Planning
This section demonstrates the effectiveness of relevance in task planning. A Fast Downward planner is employed. We compare three methods: our relevance with 0.2 as thresholds, random relevance with 0.2 as thresholds, and pure planning without relevance. For each method, we evaluate the processing time (relevance computation + planning), the timeout rate (120s), and the failure rate. Results are shown in Table II. Note that the failure rate is not applicable to pure planning, as solutions are guaranteed to be found if they exist.
Relevance dramatically improves task planning performance. Compared with Pure Planning, relevance significantly reduces the time cost by up to 79.56% and the timeout rate across all problem formulations. As the problem formulation becomes more complex, we observe a notable increase in the timeout rate and planning time for Pure Plan, highlighting its inefficiency in solving complex task-planning problems. When compared with random relevance, we observe a substantial reduction in failure rates, demonstrating that our relevance successfully and accurately determines the relevant elements of the scene. Moreover, the highest failure rate after adopting relevance in the four test cases is 0.04, which demonstrates that our relevance can effectively, accurately, and robustly reduce the task planning time.